# Reasoning Runtime Behavior of a Program with LLM: How Far Are We?

### v2

@@ Abstract
Large language models for code (i.e., code LLMs) have shown strong code understanding and generation capabilities. To evaluate the capabilities of code LLMs in various aspects, many benchmarks have been proposed (e.g., HumanEval and ClassEval). Code reasoning is one of the most essential abilities of code LLMs, but existing benchmarks for code reasoning are not sufficient. Typically, they focus on predicting the input and output of a program, ignoring the evaluation of the intermediate behavior during program execution, as well as the logical consistency (e.g., the model should not give the correct output if the prediction of execution path is wrong) when performing the reasoning. To address these problems, in this paper, we propose a framework, namely $\bm{\mathcal{R}}$Eval, for evaluating code reasoning abilities and consistency of code LLMs with program execution. We utilize existing code benchmarks and adapt them to new benchmarks within our framework. A large-scale empirical study is conducted and most LLMs show unsatisfactory performance on both Runtime Behavior Reasoning (i.e., an average accuracy of 44.4%) and Incremental Consistency Evaluation (i.e., an average IC score of 10.3). Evaluation results of current code LLMs reflect the urgent need for the community to strengthen the code reasoning capability of code LLMs. Our code, data, and $\bm{\mathcal{R}}$Eval leaderboard are available at https://r-eval.github.io.
@@ I Introduction
To provide a fair and comprehensive measure for the capabilities of code LLMs, many code-related benchmarks (e.g., HumanEval [8] and CodeXGLUE [10]) are proposed to evaluate the effectiveness of code LLMs in different tasks such as program synthesis and vulnerability detection [11]. Given that code execution provides additional information (e.g., program output) to assist with code tasks [12, 13], how to benchmark code reasoning abilities of code models with execution raises researchers’ interests [14, 15]. For example, Gu et al. [14] proposed CRUXEval to evaluate code LLMs by predicting output from input and vice versa. Typically, these works measure the model’s ability to predict and analyze the relationship between the input and output of an executable program. However, the intermediate information (e.g. execution path) during code execution is ignored, posing challenges for developers in comprehending the program’s runtime behavior.
Figure 1 shows the common concerns about the runtime behavior during program execution. Intuitively, how a program behaves under certain input can help developers better understand the code and perform debugging activities. For example, if we have concerns about the correctness of a certain statement while debugging a snippet of code, we typically first determine whether this statement is executed given the input (i.e.,  ❶ in Figure 1); If it is executed, observing the changes in variables before and after execution is a natural choice ❷; Sometimes this line of code may seem fine, so the statement immediately following it will be examined ❸; Additionally, the program output can be used to verify whether the results match the expectations ❹. Therefore, we argue that these kinds of runtime behaviors (e.g., program state and execution path) are essential for program understanding and reasoning for humans; in the meantime, it has proved to be effective for an in-depth understanding of code semantics for language models as well [16]. As previous benchmarks like CRUXEval (i.e., with ❹ and ❺) fail to evaluate if LLMs can reason about these dynamic characteristics of program, it is necessary to measure the code reasoning ability of LLMs with runtime behavior of execution. In this paper, we propose our framework, $\bm{\mathcal{R}}$Eval, to comprehensively (re)-evaluate the code reasoning ability of LLMs, which consists of two evaluation components: (i) Runtime Behavior Reasoning and (ii) Incremental Consistency Evaluation.
@@
Evaluation Component 1: Runtime Behavior Reasoning. To mitigate this limitation in previous research, we make the first attempt to systematically evaluate code LLM’s ability to reason about the runtime behavior of program execution. Specifically, we propose four evaluation tasks to achieve this goal: ❶ Code Coverage Prediction (CCP), i.e., whether a statement is executed or not; ❷ Program State Prediction (PSP), i.e., what is the value and the type of a variable; ❸ Execution Path Prediction (EPP), i.e., which is the next statement to be executed; and ❹ Output Prediction (OP), i.e., what is the output. These four tasks cover various aspects of program execution, including control flow, data flow, and type dependency, which are widely applied to prior research in software engineering such as type inference [17] and code translation [18]. Therefore, this evaluation provides a more comprehensive measure of code model’s ability to reason about executable programs in comparison with previous works.
@@
  • We construct an adapted benchmark based on HumanEval [8] and ClassEval [22] and develop an evaluation harness for our framework.
@@ II-A1 Code Execution Behavior
Previous research has leveraged code execution behavior to improve the performance of various downstream tasks, e.g., code generation [13, 12, 23] and vulnerability detection [16]. Ni et al. [12] improved code generation performance with an extra verifier, which learns the results of code execution and helps rerank generated code candidates. Chen et al. [13] utilized different kinds of feedback including output to help LLMs “self-debug” the generated code. They designed a series of prompting strategies to guide the model to refine the program automatically. In these works, pre/post-execution information like program output is applied to code generation. Furthermore, some works found the worth of dynamic features during execution and exploit them to train various language models. Liu et al. [23] pre-trained a language model to learn the execution process of program. Specifically, they represented program state as a sequence that neural models can learn from and expect the model to predict the trace. Compared to Liu et al, Ding et al. [16] proposed a pre-training technique combining both static and dynamic characteristics of the program. In summary, the aforementioned works reflect the close relationship between the behavior of code execution and the program semantics, and emphasize the importance of evaluating models for code reasoning with execution.
@@ II-A2 Code Reasoning with Large Language Models
We refer to code reasoning as predicting the code execution behavior we mentioned in the above section (i.e., pre/post-execution information and runtime information). In the task of code reasoning, the LLM needs to predict the program behavior without execution.
Recently, some works have proposed different evaluation approaches for the code reasoning abilities of code LLMs. For example, Gu et al.  [14] proposed CRUXEval which requires LLMs to reason about pre/post-execution information such as input and output. Following this study, similar to the idea of CRUXEval, Liu et al. [15] extended the evaluation tasks (i.e., predicting input and output) to the natural language specification. However, their evaluation approaches are still limited to pre/post-execution information and ignore intermediate runtime behavior. By contrast, our work goes a step further to measure how the model learns the runtime behavior during execution, which shows promising potential in helping program comprehension as said above. We notice that a recent work [24] aimed to simulate the code execution process with code LLMs. They used the analogy of a large language model to a CPU to explore the process of a program executing code, paying more attention to algorithm complexity and structure. Different from the aforementioned studies, our framework is not only limited to algorithm problems (e.g., competition-level ones), but also suitable for general programming scenarios (e.g., more real-world projects). In addition, we also explore detailed runtime behavior like code coverage and execution path, containing more runtime information.
@@ II-B Consistency for Large Language Models
Semantic Consistency. Semantic consistency refers to the same decisions on semantically equivalent texts of LLMs [25]. For example, the model should provide similar and even the same answers in the face of two meaning-perserving questions. In the realm of software engineering, this feature is generally utilized for the unsupervised evaluation for code LLMs [21, 20]: Min et al. [21] evaluated the self consistency of code LLMs by comparing the functional correctness of two code snippets: one code is generated using a human-written description and the other is generated iteratively with the summary of the previously generated. Allamanis et al. [20] introduce round-trip correctness which aligns code and NL to perform unsupervised evaluation for code LLMs. The aformentioned works leveraged the back translation between NL and PL iteratively generated by the model and conduct the semantic or functional comparison. However, they are restricted to the context of semantic consistency in the context of NL and PL.
@@ II-C1 Code LLMs
Large language models for code are LLMs specialized for the generation and understanding of PLs. For example, CodeLlama family models [6] inherit the architecture of Llama2 [30] and are further pre-trained on extra code corpora. Its three variant models (i.e., base, instruct, and python-specialized) are designed for distinct programming scenarios. StarCoder2 [31] is a series of code LLMs developed by the BigCode Project, which achieves competitive performance with other similar-sized models. These models are trained on The Stack v2 dataset [31] whose data size is four times larger than its first generation. CodeGen2.5 models [32] are improved versions of their previous models (e.g., CodeGen2) for program synthesis. It is claimed that their performance gains mainly come from optimizations such as training and sampling strategies.
Code reasoning is one of the most important capabilities of LLMs related to code [15, 14], but few works are committed to its evaluation. In this paper, we aim at comprehensively evaluating their reasoning capabilities for programming languages based on execution.
@@ II-C2 Benchmarks
Prior studies proposed a variety of code understanding tasks including code search, type inference, and code translation. Evaluating the ability of code understanding is essential for code LLMs. Lu et al. [10] propose CodeXGLUE, a comprehensive benchmark for code models supporting 10 tasks related to code and text. Cassano et al. [36] create a hand-crafted benchmark to evaluate the instruction following ability on code editing. Khan et al. [37] introduce a large-scale multilingual multitask benchmark that consists of numerous executable coding examples. Compared to the above benchmarks for code understanding, we propose a comprehensive framework for code reasoning from the perspective of runtime behavior, which provides a different point of view for the evaluation of code LLMs.
Figure 2: Overview of our framework. Benchmark Construction: we adapt base benchmarks to fit in our framework by execution. Runtime Behavior Reasoning: we propose four tasks including CCP, PSP, EPP, and OP, which challenge LLMs to perform code reasoning. Incremental Consistency Evaluation: we evaluate if the model can maintain consistency on sequentially related tasks (i.e., Incremental Consistency).
@@ III REval Framework
In this section, we first introduce the overview of our evaluation framework $\bm{\mathcal{R}}$Eval, and then describe the two evaluation components in detail, namely, Runtime Behavior Reasoning and Incremental Consistency. In the end, we describe how to construct the corresponding benchmark under our framework.
@@ III-A Overview of Framework
Figure 2 shows an overview of our framework, which aims to challenge code LLMs to reason how the program behaves during execution. To achieve this, we adopt two different perspectives for the abilities of code reasoning: (1) Runtime Behavior Reasoning; and (2) Incremental Consistency Evaluation.
For Runtime Behavior Reasoning, we focus on whether the code model can correctly predict the intermediate states of program execution given an executable program and input (as well as other contexts in the base benchmark). We select four different dimensions of runtime behavior, including code coverage, execution path, program state, and output, each of which corresponds to a specific sub-task under Runtime Behavior Reasoning. We present the task description and evaluation metric in Section III-B for each sub-task.
Besides the standalone metrics that measure a single capability of the models, we propose a novel idea named Incremental Consistency to assess the consistency across a series of incremental tasks during code reasoning. As the knowledge required to finish the latter task contains that of the former task in Runtime Behavior Reasoning, the difficulty increases progressively in order and we can utilize this characteristic to evaluate Incremental Consistency of LLMs with existing predictions (See Section III-C for details).
@@ III-B Runtime Behavior Reasoning
Runtime behavior refers to the intermediate state and information during the program execution, such as code coverage and variable values, which are widely mentioned in previous research [16, 38]. To evaluate the capability of reasoning on program runtime behavior for code models, we analyze and select four representative dimensions of intermediate information during execution, including code coverage, execution path, program state, and output prediction. Corresponding to these features, we propose four sub-tasks for Runtime Behavior Reasoning and introduce them in detail.
@@ III-B1 Code Coverage Prediction (CCP)
Code coverage measures the proportion of code covered by a test suite [39]. Recent research [38] utilized this idea to challenge the model to predict whether the statements in the program can be executed or not. Hence, we exploit the code LLM to judge if the specific statement is executed given the input of a test case.
@@
  • Evaluation Metrics. In this task, Accuracy (Acc.) presents the percentage of correct coverage predictions. For our benchmark that consists of $N$ number of $(\mathbf{P},\mathbf{X},I)$ pairs, Accuracy can be computed as:
@@ III-B2 Program State Prediction (PSP)
  • Evaluation Metrics. In this task, Accuracy measures the percentage of correct value and type predictions:
@@ III-B3 Execution Path Prediction (EPP)
In this task, we refer to the execution path of the program as an ordered sequences of statements. As the granularity of our context is statement level, we challenge the code model to predict the next statement to be executed given a specific statement. A code LLM skilled in code reasoning should be capable of telling where the control flow of the program is going and, consequently can predict the next executed statement naturally.
@@ III-B4 Output Prediction (OP)
This task is to directly generate the output of a program with the given input, which is applied in previous code reasoning work [14, 15]. To accurately predict the output of a program, code LLMs should be capable of controlling and simulating the whole execution process, which places high demands on the code reasoning ability. To evaluate the correctness of the outputs, we utilize test cases (i.e., a collection of assertion statements) in the base benchmarks. This approach is also applied in various code benchmarks [8, 22].
@@
    where $N^{\prime}$ equals to the number of different $(\mathbf{P},\mathbf{X})$ pairs in our benchmark.

### v3

@@ Abstract
Large language models for code (i.e., code LLMs) have shown strong code understanding and generation capabilities. To evaluate the capabilities of code LLMs in various aspects, many benchmarks have been proposed (e.g., HumanEval and ClassEval). Code reasoning is one of the most essential abilities of code LLMs (i.e., predicting code execution behaviors such as program output and execution path), but existing benchmarks for code reasoning are not sufficient. Typically, they focus on predicting the input and output of a program, ignoring the evaluation of the intermediate behavior during program execution, as well as the logical consistency (e.g., the model should not give the correct output if the prediction of execution path is wrong) when performing the reasoning. To address these problems, in this paper, we propose a framework, namely $\bm{\mathcal{R}}$Eval, for evaluating code reasoning abilities and consistency of code LLMs with program execution. We utilize existing code benchmarks and adapt them to new benchmarks within our framework. A large-scale empirical study is conducted and most LLMs show unsatisfactory performance on both Runtime Behavior Reasoning (i.e., an average accuracy of 44.4%) and Incremental Consistency Evaluation (i.e., an average IC score of 10.3). Evaluation results of current code LLMs reflect the urgent need for the community to strengthen the code reasoning capability of code LLMs. Our code, data and $\bm{\mathcal{R}}$Eval leaderboard are available at https://r-eval.github.io.
Index Terms:
Code Reasoning, Large Language Model, Benchmark
@@ I Introduction
To provide a fair and comprehensive measure of the capabilities of code LLMs, many code-related benchmarks (e.g., HumanEval [8] and CodeXGLUE [10]) are proposed to evaluate the effectiveness of code LLMs in different tasks such as code generation and vulnerability detection [11]. Given that “executable” is a distinct feature of code compared to natural language, and code execution provides additional information (e.g., program output) to assist with code tasks [12, 13], benchmarking code reasoning abilities of code models with execution raises researchers’ interests [14, 15]. Here, code reasoning is referred to as predicting code execution behaviors (e.g., program outputs, execution paths and possible variable values) without executing the code directly. For example, Gu et al. [14] proposed CRUXEval to evaluate code LLMs by predicting output from input and vice versa. Typically, these works measure the model’s ability to predict and analyze the relationship between the input and output of an executable program. However, the intermediate information (e.g., execution path) during code execution is ignored, posing challenges to developers in comprehending the program’s runtime behavior.
Fig. 1 shows common concerns about the runtime behavior during program execution. Intuitively, how a program behaves under certain input can help developers better understand the code and perform debugging activities. For example, if we have concerns about the correctness of a certain statement while debugging a snippet of code, we typically first determine whether this statement is executed given the input (i.e.,  ❶ in Fig. 1); If it is executed, observing the changes in variables before and after execution is a natural choice ❷; Sometimes this line of code may seem fine, so the statement immediately following it will be examined ❸; Additionally, the program output can be used to verify whether the results match the expectations ❹. Therefore, we argue that these kinds of runtime behaviors (e.g., program state and execution path) are essential for program understanding and reasoning for humans. Meanwhile, they are also proven to be effective for an in-depth understanding of code semantics for language models [16]. As previous benchmarks like CRUXEval (i.e., with ❹ and ❺) fail to evaluate whether LLMs can reason about these dynamic characteristics of a program, it is necessary to measure the code reasoning ability of LLMs with runtime behavior of execution. In this paper, we propose our framework, $\bm{\mathcal{R}}$Eval, to comprehensively (re)-evaluate the code reasoning ability of LLMs, which consists of two evaluation components: (i) Runtime Behavior Reasoning and (ii) Incremental Consistency Evaluation.
@@
Evaluation Component 1: Runtime Behavior Reasoning. To mitigate this limitation in previous research, we make the first attempt to systematically evaluate the code LLM’s ability to reason about the runtime behavior of program execution. Specifically, we propose four evaluation tasks to achieve this goal: ❶ Code Coverage Prediction (CCP), i.e., whether a statement is executed or not; ❷ Program State Prediction (PSP), i.e., what is the value and the type of a variable; ❸ Execution Path Prediction (EPP), i.e., which is the next statement to be executed; and ❹ Output Prediction (OP), i.e., what is the output. These four tasks cover various aspects of program execution, including control flow, data flow, and type dependency, which are widely applied to prior research in software engineering such as type inference [17] and code translation [18]. Therefore, this evaluation provides a more comprehensive measure of code model’s ability to reason about executable programs in comparison with previous work.
@@
  • We construct an adapted benchmark based on HumanEval [8] and ClassEval [22] and develop an evaluation harness for our framework. To facilitate further research of code reasoning, our code, data, and $\bm{\mathcal{R}}$Eval leaderboard are publicly available at https://r-eval.github.io.
@@ II-A1 Code Execution Behavior
Previous research has leveraged code execution behavior to improve the performance of various downstream tasks, e.g., program understanding [23], code generation [13, 12, 24], software testing [25, 26] and vulnerability detection [16]. Ni et al. [12] improved code generation performance with an extra verifier, which learns the results of code execution and helps rerank generated code candidates. Chen et al. [13] utilized different kinds of feedback including output to help LLMs “self-debug” the generated code. They designed a series of prompting strategies to guide the model to refine the program automatically. In these works, pre-/post-execution information, such as program output, is applied to code generation. Furthermore, some works found the worth of dynamic features during execution and exploit them to train various language models. Liu et al. [24] pre-trained a language model to learn the execution process of the program. Specifically, they represented the program state as a sequence that neural models can learn from and expect the model to predict the trace. Compared to Liu et al., Ding et al. [16] proposed a pre-training technique combining both static and dynamic characteristics of the program. In summary, the aforementioned works reflect the close relationship between the behavior of code execution and the program semantics, and emphasize the importance of evaluating models for code reasoning with execution.
@@ II-A2 Code Reasoning with Large Language Models
As introduced in Section I, in the task of code reasoning, an LLM needs to predict the program behavior without execution.
Recently, some works have proposed different evaluation approaches for the code reasoning abilities of code LLMs. For example, Gu et al. [14] proposed CRUXEval, which requires LLMs to reason about pre/post-execution information such as input and output. Following this study, similar to the idea of CRUXEval, Liu et al. [15] extended the evaluation tasks (i.e., predicting input and output) to the natural language specification. However, their evaluation approaches are still limited to pre/post-execution information and ignore intermediate runtime behavior. In contrast, our work goes a step further to measure how the model learns the runtime behavior during execution, which shows promising potential in helping program comprehension, as mentioned above. We notice that a recent work [27] aimed to simulate the code execution process with code LLMs. They used the analogy of a large language model to a CPU to explore the process of a program executing code, paying more attention to algorithm complexity and structure. Different from the aforementioned studies, our framework is not only limited to algorithm problems (e.g., competition-level ones), but also suitable for general programming scenarios (e.g., more real-world projects). In addition, we also explore detailed runtime behavior like code coverage and execution path, containing more runtime information.
@@ II-B Consistency for Large Language Models
Semantic Consistency. Semantic consistency refers to the same decisions on semantically equivalent texts of LLMs [28]. For example, the model should provide similar and even the same answers in the face of two meaning-perserving questions. In the realm of software engineering, this feature is generally utilized for the unsupervised evaluation of code LLMs [21, 20]: Min et al. [21] evaluated the self-consistency of code LLMs by comparing the functional correctness of two code snippets: one code is generated using a human-written description, and the other is generated iteratively with the summary of the previously generated. Chen et al. [29] studied the robustness of code LLMs to the variations in natural language descriptions for code generation. Allamanis et al. [20] introduced round-trip correctness which aligns code and NL to perform unsupervised evaluation for code LLMs. The aformentioned works leveraged the back translation between NL and PL iteratively generated by the model and conduct the semantic or functional comparison. However, they are restricted to the context of semantic consistency in the context of NL and PL.
@@
Figure 2: Overview of our framework. Benchmark Construction: we adapt the base benchmarks to fit our framework by execution. Runtime Behavior Reasoning: we propose four tasks including CCP, PSP, EPP, and OP, which challenge LLMs to perform code reasoning. Incremental Consistency Evaluation: we evaluate if the model can maintain consistency on sequentially related tasks (i.e., Incremental Consistency).
@@ II-C1 Code LLMs
Large language models for code are LLMs specialized for the generation and understanding of PLs. For example, CodeLlama family models [6] inherit the architecture of Llama2 [34] and are further pre-trained on extra code corpora. Its three variant models (i.e., base, instruct, and Python-specialized) are designed for different programming scenarios. StarCoder2 [35] is a series of code LLMs developed by the BigCode Project, which achieves competitive performance with other similar-sized models. These models are trained on The Stack v2 dataset [35] whose data size is four times larger than its first generation. CodeGen2.5 models [36] are improved versions of their previous models (e.g., CodeGen2) for program synthesis. It is claimed that their performance gains mainly come from optimizations such as training and sampling strategies.
Code reasoning is one of the most important capabilities of LLMs related to code [15, 14], but few work are dedicated to its evaluation. In this paper, we aim at comprehensively evaluating their reasoning capabilities for programming languages based on execution.
@@ II-C2 Benchmarks
Prior studies proposed a variety of code understanding tasks including code search, type inference, and code translation. Evaluating the ability of code understanding is essential for code LLMs. Lu et al. [10] propose CodeXGLUE, a comprehensive benchmark for code models that supports 10 tasks related to code and text. Cassano et al. [40] create a hand-crafted benchmark to evaluate the instruction following ability on code editing. Khan et al. [41] introduce a large-scale multilingual multitask benchmark that consists of numerous executable coding examples. Compared to the above benchmarks for code understanding, we propose a comprehensive framework for code reasoning from the perspective of runtime behavior, which provides a different point of view for the evaluation of code LLMs.
@@ III REval Framework
In this section, we first introduce the overview of our evaluation framework $\bm{\mathcal{R}}$Eval, and then describe the two evaluation components in detail, namely, Runtime Behavior Reasoning and Incremental Consistency Evaluation. In the end, we describe how to construct the corresponding benchmark under our framework.
@@ III-A Overview of Framework
Fig. 2 shows an overview of our framework, which aims to challenge code LLMs to reason how the program behaves during execution. To achieve this, we adopt two different perspectives for the abilities of code reasoning: (1) Runtime Behavior Reasoning; and (2) Incremental Consistency Evaluation.
For Runtime Behavior Reasoning, we focus on whether the code model can correctly predict the intermediate states of program execution given an executable program and input (as well as other contexts in the base benchmark). We select four different dimensions of runtime behavior, including code coverage, execution path, program state, and output, each of which corresponds to a specific sub-task under Runtime Behavior Reasoning. We present the task description and the evaluation metric in Section III-B for each sub-task.
Besides the standalone metrics that measure a single capability of the models, we propose a novel idea named Incremental Consistency Evaluation to assess the consistency across a series of incremental tasks during code reasoning. As the knowledge required to finish the latter task contains that of the former task in Runtime Behavior Reasoning, the difficulty increases progressively in order, and we can utilize this characteristic to evaluate Incremental Consistency of LLMs with existing predictions (see Section III-C for details).
@@ III-B Runtime Behavior Reasoning
Runtime behavior refers to the intermediate state and information during program execution, such as code coverage and variable values, which are widely mentioned in previous research [16, 42]. As shown in Fig. 2, to evaluate the reasoning ability in program runtime behavior for code models, we analyze and select four representative dimensions of intermediate information during execution, including code coverage, execution path, program state, and output prediction. Corresponding to these features, we propose four sub-tasks for Runtime Behavior Reasoning and introduce them in detail.
@@ III-B1 Code Coverage Prediction (CCP)
Code coverage measures the proportion of code covered by a test suite [43]. Recent research [42] utilized this idea to challenge the model in predicting whether the statements in the program can be executed or not. Hence, we exploit the LLM code to judge whether a specific statement is executed given the input of a test case.
@@
Evaluation Metrics. In this task, Accuracy presents the percentage of correct coverage predictions. For our benchmark that consists of $N$ number of $(\mathbf{P},\mathbf{X},I)$ pairs, Accuracy can be computed as:
@@ III-B2 Program State Prediction (PSP)
Evaluation Metrics. In this task, Accuracy (Acc.) measures the percentage of correct value and type predictions:
@@ III-B3 Execution Path Prediction (EPP)
In this task, we refer to the execution path of the program as ordered sequences of statements. As the granularity of our context is statement-level, we challenge the code model to predict the next statement to be executed given a specific statement. A code LLM skilled in code reasoning should be capable of telling where the control flow of the program is going and, consequently, can predict the next executed statement naturally.
@@ III-B4 Output Prediction (OP)
This task is to directly generate the output of a program with the given input, which is applied in previous code reasoning work [14, 15]. To accurately predict the output of a program, code LLMs should be capable of controlling and simulating the whole execution process, which places high demands on the code reasoning ability. To evaluate the correctness of the output, we utilize test cases (i.e., a collection of assertion statements) in the base benchmarks. This approach is also applied in various code benchmarks [8, 22].
@@
where $N^{\prime}$ equals the number of different $(\mathbf{P},\mathbf{X})$ pairs in our benchmark.